Price:$95:00
Sales tax rate: 4%

Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

Price:$95:00
Sales tax rate: 4%
----
Do you have a question?

Answer by jim_thompson5910(35256)

4% of 95 = 0.04*95 = 3.80

So tax alone is $3.80, add this onto the price to get the final cost (with sales tax) to be 95 + 3.80 = $98.80